Montserrat Volcano Observatory, Montserrat, West Indies

Scientific Report 08
24 January 1996

There was continued slow dome growth accompanied by frequent, very small, evenly spaced seismic events at the Soufriere Hills volcano over the past week. Conditions of excellent visibility in the later part of the week enabled close visual monitoring of the dome growing process.

The first day of the week was dominated by persistent, low amplitude, broadband tremor which was recorded only at the Gages seismometer. In the late evening of 18 January, small long-period events began to develop at a rate of one every 10 minutes or so. These continued, with occasional intermittent low amplitude broadband tremor, until 23 January. On 23 January the rate of occurrence of long-period events began to subside although events showed a slight increase in amplitude. These events were located at depths between 0.8 and 1.5 km beneath the Soufriere Hills volcano. On the evening of 23 January, very small repetitive hybrid events began to appear only on the Gages seismometer, at a rate of two per minute. By the following day, the larger long period events had almost ceased, and the very small hybrid events began to accelerate in their rate of occurrence, peaking at five events per minute. Three small volcano-tectonic earthquakes were located in the southwest of the island during the week, at depths between 1 and 3 km. Throughout the week, seismic signals were recorded that are thought to be associated with rock avalanches from the lava dome.

One regional event was recorded by the local network during the week; this was located approximately 15 km north of Montserrat. It occurred at 19:01 on 18 January and had a magnitude of 3.0.

EDM measurements continued on a daily basis through the week, with each of the four deformation triangles being measured at least once. Daily changes on all lines continue to be within the error of the method, although long-term trends still suggest a very slow swelling of the volcano. A re-occupation of the Tar River to Castle Peak EDM line showed a 2.5 cm shortening in slant distance over a period of a month; this is entirely consistent with the results obtained from the regular EDM triangle. The electronic tiltmeter at Spring Hill was removed this week and a new site is being sought. Water ingress into the equipment caused its failure. The tiltmeter at Long Ground continues not to register any deformation events.

Visual observations have been made on most days during the week. The spine reported after 14 January continued to grow significantly in the northern part of the new dome. An observation flight in poor visibility on 18 January reported that large parts of the spine had collapsed and ensuing rock falls were probably the source of ash emissions at 10:15 and 13:30 on that day. On 19 January, good visibility of the northern part of the dome confirmed that the spine had collapsed. Large blocks could be seen at the base of the rockfall talus pile, which continued to grow, and slow swelling rather than vertical growth characterised this area of activity for the remainder of the week. A new spine was noted early in the week close to the southern edge of the dome and spalling of material into the southern moat area began to build a substantial talus pile. This spine collapsed over the next few days depositing a large piece of rock into the southern moat. A new spine grew rapidly between the 20 and 21 January in a very similar location to the one that had just collapsed. This spine reached approximately 25 m in height and was about 15 m in diameter at its base. Rapid vertical growth continued until a partial collapse on 23 January which was correlated with a significant ash emission at 12:36. Thereafter, spalling of material and slow swelling and vertical growth continued in this area. Four ash emission events were reported on 24 January, all associated with major rock fall events on the lava dome. Incandescent material could be seen on the northern part of the new dome from the airport and the road to Harris through much of the week. Steaming continues from a number of places on the dome, some of which are old phreatic vents and some areas of the new dome which are now showing signs of sulphur deposition.

Light ash falls have continued to occur to the west of the volcano during the week. These are associated with large rock falls from the new dome. Observations from the helicopter and from the crater rim at Chances Peak and a new position at the rim above Farrell's have enabled good correlation between seismic signals and rock fall events. Material within the rock falls continues to be hot and heat production from the dome in general appears somewhat higher than in the past. No rocks have yet been carried outside of the crater area, although all parts of the moat are now becoming filled with debris. Most rock falls appear not to be preceded by any explosion or sudden increase in seismicity.

In summary, this week has seen the continuation of slow dome growth within English's Crater, accompanied by a long-lived swarm of small, repetitive hybrid events similar to those which have been seen on previous occasions during dome growth at Soufriere Hills volcano. Good conditions for visual observations have enabled correlations to be made between seismic signals and larger rock fall events, some of which produce small ash clouds.

Montserrat Volcano Observatory